OpenStudy (anonymous):

how do I make the >= and <= signs in latex? ("greater or equal to"...and "less then or equal to")

OpenStudy (slaaibak):

le and ge

ganeshie8 (ganeshie8):

\(\le \)

ganeshie8 (ganeshie8):

\(\ge\)

OpenStudy (anonymous):

\[\ge\]

OpenStudy (anonymous):

Yay I got now

OpenStudy (anonymous):

\[\huge \le\] = Latex Command = {le} \[\huge \ge\]= Latex Command ={ge} :D

ganeshie8 (ganeshie8):

great :) we have a group in OS for Latex practicing as well http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A)

OpenStudy (anonymous):

oh wow ...cool!

OpenStudy (experimentx):

with firefox ... right click > show math as > tex command ... copy it

OpenStudy (experimentx):

or use this as reference http://www.codecogs.com/latex/eqneditor.php